At political rally, there are 20 Republicans, 14 Democrats, and 6 Independents. It is selected at random, find the probability he or she is either a democrat or an Independent, P(D or I) = 0.5 which is also the same as 50%


1
Expert's answer
2021-12-07T12:19:03-0500

Sample Space "= 20+14+6=40"


"P(D\ or\ I)=\dfrac{14+6}{40}=\dfrac{1}{2}=0.5"

"50\%"